Brazilians are quite familiar with fueling their automobiles with ethanol, a fermentation product from sugarcane.
C2H5OH(l) +3O2(g) = 2CO2(g)+3H2O(g)
Calculate the standard molar enthalpy for the complete combustion of liquid ethanol (C2H5OH) using the standard enthalpies of formation of the reactants and products. ansswer in kj
1
Expert's answer
2019-12-03T08:25:09-0500

Hess's law state :ΔH0(reaction)="\sum" ΔHf0(products) - "\sum" ΔHf0(reactants)where  ΔHf is an enthalpy of formation, the 0 - superscript indicates standard state values.


∑ ΔHf0(reactants)= ΔHf0(C2H5OH) + ΔHf0(O2)The enthalpy of formation of a simple substance is 0."\implies" ∑ ΔHf0(reactants)= ΔHf0(C2H5OH)ΔHf0(C2H5OH)=– 277,6 kJ/mole



∑ ΔHf0(products)= 2 * ΔHf0(CO2) + 3 * ΔHf0(H2O)ΔHf0(CO2)=– 393,5 kJ/moleΔHf0(H2O)(gase)= –241.8 kJ/mole∑ ΔHf0(products) = -1512,4 kJ/moleΔH0(reaction)=-1234,8 kJ/mole



[reference (table) values ΔHf0(products) ]